OpenStudy (anonymous):

Simplify the expression 5√24 * 2√28???

OpenStudy (anonymous):

5√24 * 2√28 = (5*2) * √(24*28) = 10√672 = 10√16*√42=10*4√42=60√42

OpenStudy (anonymous):

Just use the rules for radicals, and it works out pretty easily.\[5\sqrt{24} * 2\sqrt{28}=5(2 \sqrt6)*2(2\sqrt7)=40\sqrt{42}\]

OpenStudy (anonymous):

thanks!